### v1

# Is the Top Still Spinning? Evaluating Subjectivity in Narrative Understanding

Abstract
Determining faithfulness of a claim to a source document is an important problem across many domains. This task is generally treated as a binary judgment of whether the claim is supported or unsupported in relation to the source. In many cases, though, whether a claim is supported can be ambiguous. For instance, it may depend on making inferences from given evidence, and different people can reasonably interpret the claim as either supported or unsupported based on their agreement with those inferences. Forcing binary labels upon such claims lowers the reliability of evaluation. In this work, we reframe the task to manage the subjectivity involved with factuality judgments of ambiguous claims. We introduce LLM-generated edits of summaries as a method of providing a nuanced evaluation of claims: how much does a summary need to be edited to be unambiguous? Whether a claim gets rewritten and how much it changes can be used as an automatic evaluation metric, the Ambiguity Rewrite Metric (ARM), with a much richer feedback signal than a binary judgment of faithfulness. We focus on the area of narrative summarization as it is particularly rife with ambiguity and subjective interpretation. We show that ARM produces a 21% absolute improvement in annotator agreement on claim faithfulness, indicating that subjectivity is reduced.
1 Introduction
A possible solution to the problem of factual errors in LLM-generated output lies in having a separate model or process to verify factuality (Durmus et al., 2020; Laban et al., 2022; Chen et al., 2023; Tang et al., 2024). In domains such as mathematical reasoning, verifiers like this can be used not only for evaluation, but at either training time or inference time to improve models (Zelikman et al., 2022; Wang et al., 2024). However, whether an output is factual, or whether it is entailed given some input, has been shown to be highly subjective (Pavlick & Kwiatkowski, 2019; Nie et al., 2020; Jiang & de Marneffe, 2022).This kind of subjectivity is common in tasks in the social sciences and humanities. In this work, we address narrative summarization, which plays a dual role of being a useful application of LLMs in and of itself as well as a proxy task for dealing with complex issues of subjectivity in factual judgments.
Summarizing a story is a method of capturing and distilling the key details and takeways from that narrative (Kryscinski et al., 2022; Chang et al., 2024). In this way, summarization is a vehicle for examining the understanding the summarizer has of the story. Some understandings can be clearly wrong. For example, a summary of Pride and Prejudice is wrong if it says that Mr. Darcy is a poor farmer.111He is a very wealthy member of the landed gentry. Other understandings are matters of interpretation. For example, some might summarize the end of the movie Inception as saying the top wobbles and is about to fall, indicating the main character has returned to reality,222The main character spins a top to check if he is in reality or another person’s subconscious. If it falls, he is in the real world, whereas it will just keep spinning in the subconscious. At the end, it seems he is in reality but when he spins the top, the credits roll before we see if it falls. while others would disagree with this interpretation.
Summarization work has traditionally evaluated summary faithfulness as a binary judgment of whether or not each detail in a summary is faithful to the source (Durmus et al., 2020; Fabbri et al., 2021; Min et al., 2023). If the claim says the top is about to fall though, there will be disagreement on this binary label. With narratives, sometimes claim wording is ambiguous or interpretive in a way that leads to subjective judgments of faithfulness (see example in Figure 1). Prior work has shown that in practice, it is challenging for humans to agree on this binary judgment, let alone produce a reliable automatic evaluator for narrative summarization (Subbiah et al., 2024a).
To resolve this, we first remove the assumption that there can be a universally agreed upon faithful/unfaithful label for all claims. We instead use LLM-generated rewrites as a method of evaluating claims, which we term the Ambiguity Rewrite Metric (ARM). This method produces a binary judgment (whether or not a claim is rewritten), as well as a correction of the issues in the claim. The degree to which the claim changed also provides signal on how flawed the original claim was. For example, instead of labeling the claim about the spinning top as true or false, a rewrite could specify that the top wobbles and we do not see whether or not it falls. We can quantify the amount of rewriting either through edit distance or the number of eplanation points necessary to justify the changes.
We additionally uncover the types of ambiguities in narrative claims that lead to disagreements. We distinguish between intentional ambiguity introduced by the story author and unintentional ambiguity introduced by the summary writer. We find that most ambiguities in the summary claims are unintentionally introduced by the summarizer. Finally, we further motivate this problem by demonstrating that human-written summaries also exhibit ambiguities, indicating they are an inherent part of the task. Our contributions include:
1.
An extended task definition of faithfulness in narrative summarization that allows for ambiguity in claim semantics.
2.
Human-annotated subjectivity labels and human-written summaries for the StorySumm dataset to instantiate this task.
3.
A new rewriting-based evaluation method, the Ambiguity Rewrite Metric (ARM), as an automatic evaluation for this task.
2 Background
Narrative summarization involves ambiguity and subjectivity (Subbiah et al., 2024b). Ambiguities are phrases in the story or summary that have different interpretations. For example, in the introductory example in Figure 1, the story implies an alien abduction without specifying it, leaving this interpretation ambiguous. Stories intentionally use ambiguity. Summaries can unintentionally introduce subjectivity through resolving ambiguity with explicit interpretation or using paraphrase with slightly different semantics. When evaluating the faithfulness of summaries, using only a binary faithfulness label can lead to disagreement due to the subjectivity involved (Subbiah et al., 2024a). Importantly, this disagreement is beyond annotator error as the annotators have legitimate reasons for their choices of labels.
Instead of assigning one binary label to a summary or claim, we consider the following subtasks in evaluating narrative summaries, similar to Wadhwa et al. (2024):
1.
Detecting ambiguities in a summary claim. In Figure 1, the red circle indicates the ambiguous word “aliens” in the summary claim.
2.
Fixing ambiguities in summary claims by producing rewrites of the claims. In Figure 1, the rewrite changes the wording of the abduction to match the story.
3.
Explaining what the ambiguity is with an explanation . In Figure 1, the explanation indicates the difference between the original summary claim and the story.
We hypothesize that fixing ambiguities in summary claims can reduce their subjectivity and produce more objectively faithful claims. Approaching evaluation in this way therefore focuses on measuring how far claims are from objective faithfulness rather than trying to assign an objective label up front. We therefore propose using LLM rewrites as an automatic method of evaluation, the Ambiguity Rewrite Metric (ARM). This type of evaluation has been shown useful when executed by humans but we demonstrate that LLM-generated rewrites are an effective evaluation method (Nanba & Okumura, 2004; Liu et al., 2023; yao, ).
3 Methods
We consider the setting of a narrative which is summarized into a summary . An LLM for this task places a distribution and we consider sampling predicted summaries . Each can be viewed as a collection of claims , which can be computed through a decomposition process. In our case, we consider each sentence in a summary as a claim. Instead of seeking a binary label, we consider each claim to have one of four faithfulness statuses: . N/A claims are those which just provide commentary on the story and are meant to be subjective interpretation by nature (e.g., Overall, this is a story about love and how it overcomes obstacles). We primarily focus on evaluating supported, unsupported, and ambiguous claims.
Objective claims, those which are not ambiguous, are optimal from an evaluation perspective because annotators are able to agree on their faithfulness in relation to a story. Our method leverages this insight in evaluation.
3.1 Claim rewriting
We use a rewrite model to rewrite each claim , producing a rewrite and optionally an explanation of the rewrite consisting of individual points, such that . In practice, we use an LLM as , prompting it to rewrite claims with ambiguities or unfaithful details or just repeat the original claim wording if there are no issues.
Rewrites provide three quantitative feedback signals for automatic evaluation:
1.
Whether , which provides a binary label with 1 indicating does not contain ambiguities and 0 indicating it does.
2.
The edit distance between and , which indicates how much the rewriting process changed in to address the ambiguities.
3.
The number of points discussed in the explanation, , which indicates the number of ambiguities addressed.
Rewrites additionally provide qualitative feedback: (1) the wording changes between and clearly indicate which phrasing is ambiguous in and how to correct this, and (2) the explanation explains in natural language the issues with .
3.2 Using rewrites for evaluation
We can use the benefits of rewrites discussed in Section 3.1 to address the evaluation subtasks introduced in Section 2. For detecting ambiguities, we can use the binary label to detect if a claim contains ambiguities. The qualitative feedback provided by additionally identifies what these ambiguities are. The rewrite then fixes the ambiguities in . In this way, rewrites elegantly accomplish multiple tasks in one step. Finally, the explanation explains what is ambiguous in that is addressed by .
The central question we ask is, does resolving ambiguities make claims more objectively faithful to humans? If so, rewriting claims can serve as an interpretable evaluation of where a claim lies on the ambiguity spectrum in relation to objective faithfulness. When we use as a binary label for detecting ambiguities, we refer to this as the Ambiguity Rewrite Metric (ARM). When we use the rewrites as qualitative feedback, we refer to them as just the rewrites.
4 Experimental Setup
4.1 Dataset
We use the StorySumm dataset (Subbiah et al., 2024a) to test our task definition and evaluation method. StorySumm consists of 32 English stories collected from amateur writing subreddits, and each summarized by three GPT or Claude series LLMs. In total, there are 96 summaries, consisting of 568 sentences. Each summary and sentence is labeled as faithful, unfaithful, or N/A, relative to the story. There are labels from multiple annotators along with written explanations of unfaithful labels.
Subjectivity annotation
Two of the authors of this paper assign a new label of objective or subjective for each claim in the dataset. They first read through annotator disagreements and explanations to identify a set of ambiguity types in the reasons why claims are ambiguous. They then code each claim with these types, and finally adjudicate any disagreements. Subjective claims are considered any claims assigned one of these types; the breakdown by types is discussed in Section 5.4. Objective claims can be faithful or unfaithful. In Figure 2, we see the breakdown of subjectivity labels for faithful vs. unfaithful claims. Most of the unfaithful claims are also labeled as subjective. This overlap indicates that subjectivity is a challenging part of labeling faithfulness in claims.
Establishing subjectivity in this task
To demonstrate whether subjectivity is an inherent part of this task or is introduced by LLMs in their summaries, we compare the LLM summaries against human-written summaries.333We release the human summaries and subjectivity annotations at: https://github.com/melaniesubbiah/storysumm We recruit graduate students in a computer science department to write these summaries who we trust to complete the task without LLM assistance. We collect five summaries from three students, resulting in 15 summaries, each for a different story in StorySumm, and 108 claims to evaluate.
To validate our definition of subjectivity, we also test whether inserting or removing the types of ambiguities we look for affects annotator agreement on faithfulness labels in the way we expect. As part of our subjectivity annotation, we identify four types of ambiguities (discussed in detail in Section 5.4). We generate synthetic claims using Claude-3.5 with prompts designed to produce subjective variants of objective claims in StorySumm and objective variants of subjective claims (see prompts in Appendix A.1). Each prompt introduces or corrects one of the ambiguity types we identify in our annotation.
4.2 Human studies with Upwork annotators
We perform several human studies, each using three Upwork annotators who pass a pilot screening. In each case, we show annotators a story and summary from StorySumm and ask them questions about a specific claim in the summary (see Appendix Figures 5 and 6 for the study interfaces). We use annotator disagreement on a faithfulness label for a claim as a proxy for subjectivity. While some disagreement will always arise from annotator error, significantly greater disagreement in comparing claim settings indicates more subjectivity.
We first study whether ambiguities are essential to address in evaluation:
Are ambiguities inherent to this task? For the human-written summaries, we ask whether each claim in each summary is faithful to the story. We compare the level of disagreement between the three annotators on these human-written summaries vs. on the original LLM-generated summaries for these stories. This study indicates whether humans also introduce subjective claims in narrative summarization.
Do ambiguities impact claim subjectivity? Using the synthetic claims, we create spliced summaries that are assembled by randomly selecting an objective or subjective variant for each claim in a summary. We ask whether each claim in this synthetic summary is faithful to the story and compare annotator disagreement on claims we expect to be objective vs. subjective. This study indicates whether the ambiguities we identify in claims have a measurable impact on claim subjectivity.
In early experimentation, we find that Claude-3.5-Sonnet is a strong rewrite model for this task, so we use it with the subjectivity-targeted prompt shown in Appendix A.2 for a detailed comparison to human judgments. To compare rewrites with human judgments, we use a three stage evaluation for each rewritten claim:
Does claim rewriting reduce subjectivity? We randomly show either the original or rewrite in the summary and ask whether that claim is faithful to the story. If we see greater average agreement between annotators on rewritten claims than original claims, we know rewrites make the claims more clear and objective to evaluate.
Do rewrites improve claims? We then show whichever version of the claim was not presented in the summary as an alternate and ask which version is better. We can observe whether rewrites are significantly preferred over original claims.
Are explanations of rewrites meaningful? For rewrites which annotators prefer, we parse the LLM-generated explanation for the rewrite into individual points (see prompt in Appendix A.3). We ask annotators to judge whether each point is important to their choice for why the rewrite is better. If the explanation is accurate in relation to the story and claim, the annotator can label it as IMPORTANT to their preference or NEUTRAL to their preference, and if the explanation is inaccurate, they can label it as WRONG. These annotations indicate whether explanations discuss meaningful changes in the claims.
We compare other rewrite models quantitatively but cannot perform a human study for all models. See Appendix A.4 for details on how we validate this annotation format.
4.3 Metrics
For annotator agreement, we compute the percent of claims for which all three annotators assign the same faithfulness label. For word-level edit distance between claims, we tokenize each sentence, lowercase and stem words, remove whitespace and then compute the Levenshtein distance (Navarro, 2001) using these individual words as the atomic units. We use balanced accuracy and F1-macro scores as measures of classification accuracy on imbalanced datasests for evaluating detection accuracy for subjectivity.
For evaluating explanations, we use metrics defined as follows. Let be an individual label for an explanation point, which could take the value IMPORTANT, NEUTRAL, or WRONG. Then is the set of labels corresponding to one full explanation, and is the set of corresponding to all the rewrites in the dataset. We then define % important as the macro average of the fraction of points labeled important:
and % none important as the fraction of totally not important explanations:
We can similarly calculate % wrong and % none wrong by swapping WRONG for IMPORTANT in these equations. Intuitively, % important averages across the dataset the fraction of explanation points that are labeled IMPORTANT for a rewrite. % none important is the fraction of rewrites with no explanation points labeled IMPORTANT. This difference is shown visually in Figure 3.
For statistical significance, we report a bootstrap significance test with 10,000 trials.
4.4 Baselines and rewrite models
To compare against rewrites, we use standard methods of prompting and finetuning LLMs as baselines for detecting subjective claims. We use GPT-4 (OpenAI, 2024) and Claude-3.5-Sonnet. We try zero-shot and few-shot prompting, asking whether a claim is objective. We try a self-consistency method (Wang et al., 2022) of sampling three different zero-shot CoT answers (Wei et al., 2022) for whether or not a claim is faithful with temperature 0.7. Full prompts are in Appendix A.2.
We also compare against a fine-tuned model. We use a Llama-3.1-8B-Instruct model (Grattafiori et al., 2024) and the synthetically generated data of objective and subjective claims discussed in Section 4.1. This method results in a dataset of about 2k claims that we can finetune on. We apply LoRA (Hu et al., 2022) with a rank of 64, alpha set to 64, and a dropout rate of 0.05. For hyperparameters, we use a learning rate of 5e-5, a batch size of 8, gradient accumulation over 2 steps, and train for 1 epoch.
To compare different rewrite models, we use GPT-4o and Claude-3.5-Sonnet with a standard rewriting prompt targeting subjectivity and unfaithfulness. We additionally compare this prompt against other variants targeting only subjectivity or only unfaithfulness (see Appendix A.2 for all prompts).
5 Results
Using the models and methods discussed, we establish subjectivity as part of narrative summarization and identify why claims are subjective. We demonstrate rewriting effectively detects subjectivity and corrects it in alignment with human judgments.
5.1 Does subjectivity play a role in evaluation of this task?
We compare whether human-written summaries exhibit different levels of subjectivity than LLM-written summaries. As shown in Table 1, we find that all three annotators agree on 76% of the human-written claims compared to 73% of the LLM-generated claims in StorySumm. This is not a statistically significant difference (: 0.23), indicating that ambiguous claims are an inherent challenge in evaluating narrative summaries, and are not specific to LLMs.
| % Agree | |
| Summary writer: | |
| Human | 75.93 |
| LLM | 72.74 |
| Synthetic claim type: | |
| Objective | |
| Subjective | 26.56 |
We compare whether claims generated to introduce or address the ambiguities we identify exhibit greater subjectivity or objectivity respectively. As shown in Table 1, 27% of synthetically generated subjective claims produce annotator agreement compared to 56% of synthetically generated objective claims. This difference is statistically significant (-4), demonstrating the ambiguities we identify lead to annotator disagreement. Given that these claims are synthetically generated, we do not see perfect agreement or disagreement for either category. Many of the objective claims which still produce disagreement are due to annotator error, and the subjective claims with agreement are generally due to synthetic data generation introducing an easily identifiable inconsistency.
5.2 Can rewrites detect subjectivity?
| Method | Balanced accuracy | F1-macro |
| Llama-3.1-8B finetuned | 49.62 | 0.50 |
| Claude-3.5 zero-shot | 63.85 | 0.61 |
| Claude-3.5 few-shot | 66.21 | 0.62 |
| Claude-3.5 self-consistency | 50.42 | 0.50 |
| Claude-3.5 ARM (ours) | 69.15 | 0.53 |
| GPT-4 zero-shot | 58.04 | 0.57 |
| GPT-4 few-shot | 58.94 | 0.56 |
| GPT-4 self-consistency | 57.73 | 0.59 |
| GPT-4 ARM (ours) | 63.55 | 0.58 |
The ARM metric labels claims which get rewritten as subjective and labels them objective otherwise. We compare how well ARM detects subjectivity relative to the baselines detailed in Section 4.4 in Table 2. We find that finetuning a detection model on the synthetic dataset does not transfer effectively to the original claims. In zero-shot and few-shot settings, we see that Claude is overall a stronger model for detecting subjectivity than GPT-4. Self-consistency does not work well for Claude but has comparable performance to other methods for GPT-4. ARM is a strong method for detecting subjectivity across both models, and the strongest when looking at balanced accuracy. Overall, ARM is a stronger or comparable method to other baselines for binary evaluation while providing the added benefits of the rewrite itself and explanation. We report additional results with alternate prompts and against faithfulness labels in Appendix A.7 but observe that ARM is fairly consistent across different prompts.
5.3 Do rewrites align with human judgments of faithfulness?
| Metric | Original claims | Rewrites |
| Agreement | 36.36 | 57.45* |
| Faithful | 20.45 | 89.36** |
| Preferred | 23.08 | 76.92** |
We take the subjectivity-focused rewrites using Claude (prompt in Appendix A.2) and test whether its edits align with human judgments and preferences. In Table 3, we see that rewrites exhibit statistically significant gains over original claims in annotator agreement and faithfulness. When human judges are asked whether they prefer the original claim versus the rewritten claim, they prefer the rewrites 77% of the time.
In Table 4, we report the quality of explanations. We find that on average 69% of explanation points are labeled as IMPORTANT by majority vote. 99% of explanations have no points labeled WRONG by majority vote. Finally, only 5% of explanations have no points considered important for an individual annotator. These percentages show that the vast majority of explanations and changes made in the rewrites are either important to fix or neutral. We note that the explanation parse often repeats the same point multiple times (see example explanations in Figure 3), which is why we consider both % important and % wrong.
These numbers indicate that rewrites significantly improve claims in objectivity and faithfulness and using them in conjunction with their explanations is a meaningful evaluation signal.
| Important | Wrong | |||
| Method | % Important | % None important | % Wrong | % None wrong |
| Individual | 84.11 | 5.29 | 7.52 | 87.83 |
| Maj. vote | 68.72 | 24.39 | 0.56 | 98.78 |
5.4 Qualitative analysis of ambiguities
We uncover specific reasons why claims are ambiguous and analyze whether there are specific types of ambiguities that are harder for rewrites to detect. Two of the authors of this paper perform inductive thematic analysis (Bowman et al., 2023) with adjudication to identify a taxonomy for why claims are subjective (see Appendix A.6 for more details). Using this taxonomy, they revisit each subjective claim and label it with an ambiguity code. The taxonomy identifies four ambiguity types:
1.
Wording: A word or phrase is used in the summary which has overlapping meaning with the wording used in the story but also lends itself to other interpretations. Depending on someone’s interpretation, the two meanings may not fully overlap.
2.
Detail: A very minor detail is assumed in the summary which is not explicitly stated in the story. Many people find this to be a reasonable assumption and therefore faithful, while others view it as unfaithful.
3.
Causation: The summary skips important causal details for an event. Some people find this to be a reasonable abbreviation while others feel it fundamentally changes one’s understanding of what happened.
4.
Explicit: The summary makes explicit details that are intentionally left ambiguous or only implied in the story. Some people like this interpretive jump while others feel it misrepresents the nature of the story.
In Figure 4, we see the most common ambiguity type in the LLM-generated subjective claims in StorySumm is Type 1, meaning many claims use vague or misleading wording. This result is consistent with prior work (Subbiah et al., 2024b; Kim et al., 2024). Type 4 is the smallest category, indicating that most ambiguities are unintentionally introduced by the summary writer (Types 1-3). We also see that on average, there does not seem to be one type of ambiguity that is detected substantially more or less by the LLM rewrite metrics. We compare ARM’s recall across the different ambiguity types but do not find that one type of claim is missed more often than others (see Appendix Figure 7).
6 Related Work
Ambiguity in evaluation Ambiguity has been studied in natural language entailment (Pavlick & Kwiatkowski, 2019; Nie et al., 2020; Jiang & de Marneffe, 2022), a related task to summary faithfulness evaluation. Other work has studied how to improve or manage annotator disagreement (Uma et al., 2022; Krishna et al., 2023; Min et al., 2023). Most similar to ours is Koupaee et al. (2025), which examines ambiguity in evaluating meeting summaries. Their taxonomy of ambiguities supports our findings and by working with narratives, we incorporate the added challenge of intentional ambiguity in the source text.
Edits as feedback Recent work has explored the use of natural language feedback to improve model outputs in different settings, such as math reasoning (Madaan et al., 2023; Xu et al., 2024b) and summarization (Liu et al., 2023; Zhang et al., 2023). This feedback can be human feedback (Madaan et al., 2023) or automatic feedback from models (Gao et al., 2023; Ye et al., 2023; Wadhwa et al., 2024; Xu et al., 2024a). While most of these methods follow a critique-and-refine framework to improve models’ outputs, we propose a rewriting-based evaluation method that leverages edits themselves as the feedback signal.
Evaluating narrative summaries Prior work has introduced datasets and methods for evaluating narrative summaries. Early work provided reference summaries from online study guide and TV episode synopsis websites to enable reference-based evaluation metrics (Ladhak et al., 2020; Kryscinski et al., 2022; Chen et al., 2022). As LLMs have trained on more and more online data, more recent work has evaluated LLMs as evaluators on recently published books or unpublished work (Chang et al., 2024; Kim et al., 2024; Subbiah et al., 2024b; Karpinska et al., 2024).
7 Conclusion
In this work, we expand the considerations for evaluating faithfulness in narrative summarization to include ambiguity and subjectivity. We propose using rewrites as automatic evaluation for summary claims and demonstrate their efectiveness on the StorySumm dataset. We release additional labels for StorySumm for claim subjectivity and ambiguity types. In the future, we hope this evaluation methodology can be tested on other challenging evaluation tasks in the humanities. We believe rewriting-based evaluation could also be used in RL training for improving reasoning about implicit meaning in the humanities.
Ethics Statement
We follow protocol approved by our IRB for the human annotation work in this study. One of the authors holds an equity interest in OpenAI.
Appendix A Appendix
A.1 Synthetic data generation prompts
The following prompts are filled with a story and summary or story and claim pair in the %s fields.
A.2 Baseline and rewrite prompts
The following prompts are filled with a story, summary, and claim triple in the %s fields.
A.3 Explanation parsing prompt
A.4 Annotation format validation
We check for confounding factors in the task format for the human annotations of rewrites in several ways. First, we randomize whether the rewrite is shown to the annotator in the summary or as the “alternate”. We find there is not a significant difference (: .08) in which claim is preferred based on which position it is shown to annotators in (57.1% for in-summary vs. 42.9% for alternate). These numbers indicate annotators are not biased to prefer whichever claim is shown to them as the “original” vs. “alternate”.
Additionally, we include three decoy explanations that do not make sense and annotators should reject to check that annotators are not just convinced by any explanation. The decoy explanations are marked as WRONG in 5/6 instances indicating that annotators are not just convinced by any explanation. We remove results on these decoy explanations from the results discussed in the paper.
A.5 Human annotation interfaces
A.6 Additional ambiguity analysis details
For the inductive thematic analysis, the two authors involved write an explanation for the ambiguities involved in each subjective claim. They then discuss these explanations and arrive at four types of ambiguities. They go back and code each subjective claim with one of these types and discuss and adjudicate any disagreements.
They additionally label a small number of claims with a type 5 not discussed in the main text. Type 5 indicates the story is too confusing to determine its intent but the summary sentence itself is written clearly. We do not include type 5 claims in the analysis as they result from unintentional ambiguity introduced by the story writer which is beyond the scope of this paper.
While the writer of a story may intentionally use ambiguity in their story which leads to subjective viewpoints on its meaning, a summary writer should not intentionally introduce ambiguity. Types 1-3 capture types of ambiguities in summary claims that are unintentional on the part of the summary writer. Type 4 deals with intentional ambiguity by the author. We observe that type 4 ambiguities may be okay in summaries that should interpret the story as well as summarizing it, while types 1-3 are always undesirable as they obscure meaning.
In Figure 7, we show the recall of Claude and GPT-4 rewrite metrics averaged across the three different rewriting prompts for each ambiguity type. We do not observe a significant difference in recall for the different types.
A.7 Does the rewrite model and prompt affect results?
| Method | Subj. | Unfaith. | Subj. unfaith. | # Rewrites | Avg. edit dist. |
| Claude subj. | 67.89 | 67.42 | 69.00 | 246 | 0.40 |
| Claude both | 69.15 | 68.05 | 69.96 | 341 | 0.61 |
| Claude inconsist. | 66.7 | 68.21 | 63.56 | 218 | 0.46 |
| GPT-4 subj. | 53.95 | 55.42 | 54.61 | 486 | 0.17 |
| GPT-4 both | 63.55 | 64.95 | 63.39 | 203 | 0.22 |
| GPT-4 inconsist. | 64.01 | 63.09 | 63.56 | 390 | 0.21 |
We assess how sensitive rewrites are to prompting method and model choice in Table 5. We evaluate detection against subjectivity labels, faithfulness labels, and claims that are subjective or unfaithful. We use the rewriting prompts shown in Appendix A.2. We once see that Claude is a stronger model for rewriting in terms of binary classification of subjectivity or faithfulness. We see that GPT-4 provides the benefit of making more minimal changes to claims though as its rewrites have less than half the edit distance relative to Claude. Overall, the LLM used for rewriting matters more than the prompt with the exception of the subjectivity-targeted prompt for GPT-4. This setting produces the most rewrites by far which leads to its poor performance.